### v1

# Fine-Tuning Universal Machine-Learned Interatomic Potentials: A Tutorial on Methods and Applications

Abstract
Universal machine-learned interatomic potentials (U-MLIPs) have demonstrated broad applicability across diverse atomistic systems but often require fine-tuning to achieve task-specific accuracy. While the number of available U-MLIPs and their fine-tuning applications is rapidly expanding, there remains a lack of systematic guidance on how to effectively fine-tune these models. This tutorial provides a comprehensive, step-by-step guide to fine-tuning U-MLIPs for computational materials modeling. Using the recently released MACE-MP-0 as a representative example, we cover key aspects of the fine-tuning process, including dataset preparation, hyperparameter selection, model training, and validation. The effectiveness of fine-tuning is demonstrated through case studies involving force prediction in solid-state electrolytes, stacking fault defects in metals, and solid–liquid interfacial interactions in low-dimensional systems. To support practical applications, we include code examples that enable researchers, particularly those new to the field, to efficiently incorporate fine-tuned U-MLIPs into their workflows.
preprint: APS/123-QED
I INTRODUCTION
Machine-learned interatomic potentials (MLIPs) have become an essential tool in materials modeling, offering near-density functional theory (DFT) accuracy at significantly reduced computational cost [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. A broad range of MLIP frameworks have been developed in recent years, including neural network potentials [1, 4], kernel-based approaches [2, 5], and equivariant graph neural networks [7, 6], each demonstrating excellent performance across diverse materials systems. These models enable accurate and large-scale atomistic simulations, and their integration with simulation platforms such as ASE [15] and LAMMPS [16] has facilitated their widespread adoption in the field. For comprehensive reviews on MLIPs, the reader is referred to Refs. [17, 18, 19, 20, 21].
Traditionally, most MLIPs are trained on narrowly defined chemical or structural domains tailored to specific applications. While such models can achieve high accuracy within their training scope, they often lack transferability to systems beyond that domain. In response, there has been growing interest in a new class of models, often referred to as general-purpose, foundation, or universal MLIPs (U-MLIPs) [22, 23, 24, 25, 26, 27]. Inspired by foundation models in natural language processing and computer vision [28], these models are pre-trained on large and diverse materials datasets [29, 30, 31], aiming to provide broad coverage across the periodic table and structural configuration space. They offer strong baseline performance and could be readily applied to new systems with minimal additional training (i.e., fine-tuning), substantially lowering the barrier to entry for MLIP applications in materials science. In Table 1, we summarize representative U-MLIPs, along with their publication year, underlying architectures, and training data characteristics.
Table 1: List of the representative U-MLIPs. See Appendix V.1 for further details.
Model Name Year Architecture Data Description
M3GNET [27] 2022 SchNet [32] Materials Project (MP), 89 elements, 62783 compounds
CHGNet [23] 2023 GNN + Charge MP + Trajectory, 89 elements, 146000 compounds
ALIGNN-FF [26] 2023 Line GNN JARVIS-DFT, 72708 compounds
PFP (Matlantis) [33] 2023 Tensorial GNN Custom, $\approx$10 million configurations
GNoME [24] 2023 NequIP [8] MP + Custom, $\approx$89 million configurations
DPA-1 [25], DPA-2 [25] 2023, 2024 DeepMD [4] Open-access datasets including Alloy, OC2M, and others.
MACE-MP-0 [22] 2024 MACE [6] Same data as used to build CHGNet
SevenNet-0 [34] 2024 NequIP [8] Same data as used to build CHGNet
MatterSim [35] 2024 Graph transformer MP+Alexandria+Custom, $\approx$17 million configurations
EquiformerV2-OMAT24 [31] 2024 EquiformerV2 [36] MP+Alexandria+Custom, $\approx$118 million configurations
Among existing U-MLIPs, MACE-MP-0 stands out as a near state-of-the-art model, demonstrating good accuracy and robust stability in molecular dynamics simulations across a wide range of applications [22]. It serves as the central focus of this work. MACE-MP-0 is built upon the MACE architecture [6], which integrates the atomic cluster expansion (ACE) formalism [5, 37] with efficient tensor decomposition techniques [38] and higher-order equivariant message passing. Leveraging this design, MACE-MP-0 is pre-trained on a large dataset from the Materials Project, and achieves strong performance across diverse materials benchmarks. Several subsequent variants of the model have been introduced recently to address specific issues such as repulsion, phonons, and high-pressure stability, further broadening its applicability.
While MACE-MP-0 and its varaints have demonstrated robust performance in molecular dynamics simulations [22], its predictive accuracy—like that of other U-MLIPs—remains limited for certain tasks, particularly in modeling mechanical properties such as elastic constants and stacking fault energies in elemental alloys [39]. This underscores the need for fine-tuning, which adapts pre-trained models to specific systems by refining their parameters using high-fidelity computational or experimental data. Although the fine-tuning of atomistic foundation models has garnered growing attention, comprehensive and systematic studies remain limited, especially practical tutorials that guide beginners through the fine-tuning process. Recent works have started addressing this gap [40, 41, 42, 43, 44, 45, 46, 47, 48], and continued efforts are expected to clarify best practices across various application scenarios.
Beyond the pursuit of higher accuracy, a central motivation for fine-tuning in the context of machine-learned interatomic potentials lies in a fundamental question: given a fixed dataset, should one fine-tune an existing foundation model or train a new model from scratch? In practice, fine-tuning is often more efficient in terms of training time and data usage, and in many cases yields comparable or even superior accuracy. While foundation models, pre-trained on large and diverse datasets, effectively capture general chemical and structural patterns, they often lack the precision required for specific systems. Fine-tuning addresses this limitation by incorporating system-specific data, thereby enhancing accuracy with significantly lower data and computational costs. Moreover, fine-tuned models typically converge faster due to the advantage of informed initialization. Therefore, when a suitable foundation model is available, fine-tuning presents a compelling alternative to training-from-scratch.
This tutorial offers a practical guide to fine-tuning U-MLIPs, with a particular focus on helping researchers who are new to the area. The goal is to provide step-by-step instructions for preparing data, training models, selecting hyperparameters, and applying the resulting potentials in various atomistic simulations. To keep the presentation accessible, theoretical details are kept to a minimum, with references provided for readers seeking a deeper mathematical and physical understanding. The methods introduced here are demonstrated using widely used tools such as ASE, LAMMPS, and the Random Batch Molecular Dynamics (RBMD) interface [49]. MACE-MP-0 is selected as the main example due to its broad applicability, high accuracy across benchmark tasks, and demonstrated stability in molecular dynamics simulations [22]. While this tutorial is based on MACE-MP-0, the fine-tuning workflow and practical considerations apply more generally to other U-MLIPs as well  [^1] A note of caution: The procedures and recommendations in this tutorial reflect the state of the field at the time of writing. Given the rapid pace of developments in model architectures, datasets, and fine-tuning strategies, updates are likely in the near future..
It is worth noting that the fine-tuning of U-MLIPs can be efficiently carried out using the RBMD package, which enables large-scale particle simulations at the nano- and microscale. In contrast to conventional molecular dynamics frameworks, RBMD leverages random batch algorithms [51, 52, 53, 54] to efficiently handle nonbonded interactions, supporting simulations of up to 10 million particles on a single GPU with a CPU core [49]. Moving forward, we aim to integrate the RBMD platform with MACE-based MLIPs for more practical and large-scale simulations. To support this integration, an open-access platform is available at https://www.randbatch.com/rbmd, where further examples and implementation scripts for realistic simulation scenarios will be released in the future.
The remainder of this tutorial is organized as follows. In Section II, we introduce the overall fine-tuning workflow for U-MLIPs, including a brief overview of the MACE-MP-0 framework, data preparation, potential training, and model validation. Section III presents three illustrative applications of fine-tuning U-MLIPs: (i) force accuracy evaluation in the solid-state electrolyte Li10GeP2S12 (LGPS), (ii) stacking fault energetics in body-centered cubic molybdenum (Mo), and (iii) surface interactions at the graphene–water interface. Finally, Section IV provides a summary and outlook on the development and fine-tuning of U-MLIP models.
II METHODS
In this section, we introduce the overall procedure for fine-tuning U-MLIPs, using MACE-MP-0 as a representative model. An overview of the fine-tuning workflow is shown in Figure 1. The section is organized as follows: Section II.1 briefly describes the MACE-MP-0 model, common fine-tuning practices, and recent implementation updates. Dataset construction, a critical component of MLIP development, is discussed in Section II.2. Section II.3 outlines the fine-tuning process, including hyperparameter choices. Model validation is addressed in Section II.4, which is a prerequisite for reliable molecular dynamics simulations.
The fine-tuning procedure is supported by the official MACE repository: https://github.com/ACEsuit/mace. Installation instructions are also available in the tutorial repository: https://github.com/John2021-hub/mace-ft-tutorial.git. For completeness, Table 8 in Appendix V.3 provides an overview of relevant repositories and their functionalities. For instance, the RBMD interface [49] supports both fine-tuning and molecular dynamics simulations with MACE-based MLIPs.
Figure 1: Overview of the fine-tuning workflow
II.1 The MACE-MP-0 Model: Architecture, Fine-Tuning, and Recent Developments
We begin by reviewing the MACE-MP-0 model in Section II.1.1. The fine-tuning procedures introduced here are broadly applicable to other U-MLIPs with only minor adjustments. For completeness, a brief overview of other representative U-MLIP models is provided in Appendix V.1.
II.1.1 Overview of the MACE-MP-0 Model
Before introducing U-MLIPs, we briefly recall the concept of MLIPs. These models map atomic positions and chemical elements to the potential energy of a given atomic system. The MACE architecture [6] is a representative example, extending the Atomic Cluster Expansion (ACE) framework [55, 5, 37], which was originally developed for MLIPs and has since found applications beyond atomistic modeling [56, 57, 58, 59]. MACE is an equivariant message-passing graph tensor network [60, 61], where many-body atomic geometry is encoded layer by layer. Messages are built from linear combinations of tensor product bases derived from two-body permutation-invariant polynomials in a spherical basis [5, 37]. Equivariance is maintained through tensor contractions with irreducible representations and generalized Clebsch–Gordan coefficients [62, 38]. The final output represents the per-atom contribution to the total potential energy. Further architectural details are provided in Appendix V.2 and in Ref. [6].
The foundation model MACE-MP-0, trained on the Materials Project dataset [63], has demonstrated strong performance in a range of applications [22], particularly in enabling stable molecular dynamics simulations across diverse chemical environments. It is released in three variants, differentiated by increasing levels of message-passing equivariance ($L=0,1,2$). A subsequent revision, MACE-MP-0b, introduced minor improvements such as refined pairwise repulsion and improved treatment of isolated atoms. Further variants have been developed to address specific limitations: MACE-MP-0b2 enhances stability under high pressure, MACE-MP-0b3 corrects phonon behavior, and MACE-MPA-0 and MACE-OMAT-0 extend training coverage. All models are publicly available at https://github.com/ACEsuit/mace-mp/. In this work, we adopt MACE-MP-0b3 as the base model for fine-tuning. For simplicity, we continue to refer to it as MACE-MP-0 unless otherwise stated.
Despite the general robustness of these models, stable MD simulations do not guarantee accurate predictions of physical properties. Fine-tuning remains essential for adapting U-MLIPs to specific systems. However, to the best of our knowledge, there is currently no systematic tutorial guiding users through the fine-tuning of U-MLIPs for practical applications. Addressing this gap serves as the central motivation for the present work.
Fine-tuning pre-trained models is widely adopted to improve accuracy and transferability in various domains, including image analysis [64, 65] and natural language processing [66, 67, 68]. In molecular modeling, fine-tuning typically follows a predictor–corrector paradigm: a U-MLIP serves as an initial predictor that provides a robust baseline for atomistic simulations. The model is then fine-tuned on a system-specific dataset, allowing it to correct and adapt to the target system with improved accuracy.
Multi-head fine-tuning extends this idea by attaching multiple task-specific output heads to a shared pre-trained backbone [69]. Each head is tailored to a different target system or property, while the backbone captures general representations. By partially freezing the backbone, this approach enables efficient training across multiple tasks and reduces redundancy. An implementation of this method is available at https://github.com/ACEsuit/mace/tree/main.
II.1.2 Recent Updates
This section briefly introduces two recent updates relevant to U-MLIP deployment: the fast equivariant kernel cuEquivariance, and an implementation of long-range dispersion corrections with efficient neighbor list handling. These tools are integrated into our examples (cf. Section III) to demonstrate the fine-tuning and application of the MACE-MP-0 model in practical simulations.
The fast equivariant kernel cuEquivariance.
Equivariance, referring to the preservation of pre-defined physical symmetry under group operations such as rotations and translations, is a key property of physically meaningful interatomic potentials. Models that are naturally equivariant are typically more data-efficient, posses better physicality and exhibit improved generalization.
cuEquivariance is an NVIDIA-developed library for building high-performance equivariant neural networks based on segmented tensor products within easily accessible Python interface. It provides a flexible API and optimized CUDA kernels, with bindings for both PyTorch and JAX. The code is available at https://github.com/NVIDIA/cuEquivariance.
Fine-tuned MACE-MP-0 models (cf. Section II.3) can be integrated with cuEquivariance in Python-based simulation environments such as ASE [15], enabling significantly faster inference—typically by a factor of 3 to 10. Example usage and performance benchmarks are provided in Section III.1. The following code snippet illustrates how to activate the fast kernel via a simple keyword:
1from mace.calculators import MACECalculator
2mace_calc = MACECalculator(model_paths="MACE.model", enable_cueq=True, device="cuda")
Listing 1: Enable cuEquivariance kernel
As reported by developers, the jax implementation is expected to have at least extra few factors of speed up further.
Long-range dispersion corrections with efficient neighbor list handling.
Frequent neighbor list updates can introduce significant computational overhead, particularly when large cutoffs are used for dispersion corrections such as DFT-D3 [70]. To mitigate this, a simple yet effective neighbor list implementation with a skin layer is provided in ASE. This approach reduces the frequency of neighbor list reconstruction and improves overall computational efficiency.
The implementation allow similar update strategy in LAMMPS and enables the reuse of neighbor lists across multiple simulation steps. An example configuration is shown below:
1calc = TorchDFTD3Calculator(atoms=atoms, device="cuda", damping="bj",
2 every=2, delay=10, check=True, skin=2.0)
Listing 2: Enable neighbor list reuse
In this example, the parameters every, delay, check, and skin control the update frequency and reuse behavior of the neighbor list:
  • every: interval between neighbor list updates.
  • delay: buffer before enforcing the next update.
  • check: whether to validate the reuse condition.
  • skin: width of the buffer zone (in Å), allowing small displacements without triggering a rebuild.
The implementation is available at: https://github.com/CheukHinHoJerry/torch-dftd.
II.2 Data Preparation
The quality of a MLIP is closely tied to the quality of the reference dataset. MLIPs are typically trained via supervised regression on a large set of atomic configurations, each labeled with total energies, atomic forces, and often stress tensors, usually computed from DFT. The same principles apply to fine-tuning U-MLIPs, where a well-designed dataset is crucial for achieving improved accuracy and transferability [71, 40, 42, 43, 72].
A key question in the fine-tuning of U-MLIPs is how to select a dataset that leads to better performance—both in terms of accuracy and data efficiency—compared to using the original foundation model or training a new model from scratch. In this section, we summarize a range of strategies, from empirical approaches based on physical intuition (Section II.2.1) to more automated data selection guided by uncertainty quantification (Section II.2.2).
II.2.1 Empirical Dataset Construction
Empirical dataset construction relies on physically motivated or experience-driven methods. These include using existing public datasets, generating configurations via random perturbations, or sampling configurations from molecular dynamics (MD) or Monte Carlo (MC) simulations. Each method offers distinct advantages and is best suited for particular types of systems or user expertise. In the following, we briefly describe each approach and offer practical recommendations based on our experience.
Using publicly available datasets.
Depending on observables or quantities of target system, open-access datasets that have been developed for machine-learned interatomic potentials can be readily used for fine-tuning U-MLIPs [29, 30, 31, 73, 74, 75]. A key consideration is format compatibility. For example, the MACE-MP-0 model expects data in the .xyz format, whereas datasets for Deep Potential (DP) models [4] are typically provided in .npy format. To address this, we provide a convenient script that converts .npy files into .xyz files:
1python transfer_npy_xyz.py train.npy train.xyz
Listing 3: Convert .npy data to .xyz
Another consideration is dataset size. Public datasets designed for training-from-scratch are often much larger than what is needed for fine-tuning. Since U-MLIPs already encode broad chemical and structural trends, effective fine-tuning can usually be achieved with significantly fewer configurations. This observation will be demonstrated in a case study in Section III.1. We note that using publicly available datasets is especially recommended for beginners, as it requires minimal setup and provides immediate access to high-quality reference data.
Random perturbation of structures.
Another common approach is to apply random displacements to atoms in relaxed structures to sample configurations near local energy minima (local equilibrium). This method is simple and widely used, especially for systems with crystalline structures. If stress data are required, small random deformations of the simulation cell (i.e., strain perturbations) should also be applied. Random perturbations can be generated using ASE’s rattle function, typically with Gaussian-distributed displacements of 0.1–1.0 Å. Care must be taken to avoid unphysical structures with atoms too close together. This issue can be mitigated by applying a Monte Carlo filtering procedure that enforces a minimum interatomic distance, where configurations with excessively short atomic separations are accepted with low probability [76]. This method is recommended for users who are familiar with atomistic simulation workflows and have sufficient computational resources. It is particularly well suited for advanced users with prior knowledge of the target system. Notably, when combined with the filtering techniques discussed later, well-designed perturbation strategies can produce compact and physically meaningful datasets that are particularly effective for fine-tuning U-MLIPs [77].
Sampling from MD or MC simulations.
When no well-defined structural prototype exists, as in amorphous phases or liquids, perturbation methods are not suitable. Instead, ab initio molecular dynamics [78, 79] or Monte Carlo simulations provide a more general approach for generating configurations. Snapshots are extracted at regular intervals along the trajectory to build the dataset. Although computationally more demanding, this approach is essential when structural templates are unavailable. To further improve dataset diversity and relevance for fine-tuning, perturbations are often applied to MD-generated structures, and in some cases, MD simulations with enhanced sampling are employed [80, 81].
II.2.2 Uncertainty-Based Dataset Construction
Manual dataset design often depends on expert intuition, which may introduce inductive biases, redundant structures, or sampling imbalances. Such issues can undermine data efficiency and hinder the generalizability of trained MLIPs. As a remedy, uncertainty-guided data selection has emerged as a systematic and efficient alternative for both training MLIPs and fine-tuning U-MLIPs [82, 83, 84, 85, 86, 87].
Uncertainty quantification (UQ) estimates the confidence of a model’s predictions and can be used to identify regions in configuration space where the model is unreliable. Incorporating such estimators enables active learning strategies that prioritize configurations with high predictive uncertainty, thereby improving accuracy while minimizing data redundancy.
Several definitions and methodologies for uncertainty estimation are commonly employed in the context of MLIPs. These approaches can be broadly categorized into frequentist and Bayesian frameworks, as summarized in Table 2. Frequentist techniques include model ensembles [88] and feature-space distance metrics [89], while Bayesian methods rely on posterior inference [85], dropout-based approximations [86], or analytical forms available in tractable models such as Gaussian processes and linear regressions [77]. In addition, hybrid approaches have been developed to separately quantify aleatoric (data-inherent) and epistemic (model-related) uncertainties [90]. A comprehensive discussion of UQ in MLIPs is beyond the scope of this tutorial. However, we emphasize that this remains an active area of research, and we refer interested readers to recent reviews for further insight [91, 92, 93].
Table 2: Representative uncertainty quantification (UQ) methods used in MLIPs.
Method Key Idea Category
Ensemble [94] Use variance across multiple independently trained models Frequentist
Dropout UQ [95] Dropout at inference to sample pseudo-ensembles Bayesian
Bayesian NN [96] Infer posterior over weights via VI [97] or MCMC [98] Bayesian
Evidential UQ [99] Predict mean and variance directly as outputs Evidential
GPR variance [100] Analytical uncertainty from Gaussian processes Bayesian
Feature distance [101] Distance to nearest training point in feature space Geometric
Aleatoric + Epistemic [102] Explicitly model data and model uncertainty Combined
Filtering a candidate set.
In cases where a large pool of configurations is available (e.g., from long MD simulations or existing datasets), uncertainty-based filtering can be applied to construct an efficient fine-tuning set. Here, an uncertainty estimator (such as an ensemble or a surrogate model with Bayesian regression) is used to assign a confidence value to each configuration. Configurations with higher estimated uncertainty are prioritized, as they are more likely to contribute new information to the model. This strategy has been used successfully in recent works, such as Ref. [77], where a linear ACE model is trained on top of an existing representation to guide uncertainty filtering.
Active learning strategy.
Active learning provides a fully iterative framework for dataset construction. Starting from a small initial set of labeled configurations, the MLIP is trained and then used to run MD simulations. Configurations from these simulations are evaluated by a UQ metric, and the most uncertain ones are selected for DFT labeling. The newly labeled data are added to the training set, and the process is repeated until convergence or a stopping criterion is met (e.g., error saturation or maximum data budget). This strategy is particularly effective for exploring diverse chemical spaces or unknown physical regimes.
While uncertainty-driven methods offer significant advantages, they are not universally applicable. Analytical UQ is often limited to linear models or kernel-based methods such as Gaussian process regression. For large, non-linear models like deep neural networks, approximate UQ methods (e.g., ensembles, dropout) may require careful calibration [103, 104]. Additionally, active learning workflows introduce algorithmic and software complexity, which may be a barrier for new users.
II.3 Fine-Tuning
This section outlines the main steps and considerations for fine-tuning the MACE-MP-0 model. Fine-tuning refers to updating the parameters of a pre-trained U-MLIP on a task-specific dataset, while keeping the overall model architecture fixed. This strategy leverages the broad physical knowledge encoded in the foundation model and enables efficient adaptation to new systems with relatively limited data.
A typical fine-tuning workflow involves constructing the training dataset, (and usually a validation dataset) set with the above mentioned dataset construction methods. Table 3 summarizes representative hyperparameters and their recommended ranges, consistent with the example script provided in Listing 4.
Table 3: Representative hyperparameters for fine-tuning MACE-MP-0. The recommended values are representative for the cases considered in this work, and may require adjustment for different systems.
Parameter Description Recommended Range
energy_weight Energy loss weight 1.0 – 20.0
forces_weight Force loss weight 1.0 – 20.0
stress_weight Stress loss weight 0.0 – 20.0
lr Learning rate $10^{-4}$$10^{-3}$
batch_size Batch size 2 – 20
ema_decay EMA decay rate 0.98 – 0.99999
swa Enable stochastic weight averaging True / False
swa_lr Learning rate for SWA phase $10^{-5}$$10^{-4}$
Listing 4 presents a fine-tuning script using the run_train.py utility provided by the MACE package:
1python3 \$MACE_DIR/mace/cli/run_train.py \
2 --name=\$RES_DIR \
3 --foundation_model="XXX.model" \
4 --model_dir=results/\$RES_DIR \
5 --log_dir=results/\$RES_DIR \
6 --checkpoints_dir=results/\$RES_DIR \
7 --results_dir=results/\$RES_DIR \
8 --train_file="train.xyz" \
9 --valid_file="valid.xyz" \
10 --energy_weight=1.0 \
11 --forces_weight=10.0 \
12 --stress_weight=1.0 \
13 --loss="universal" \
14 --forces_key=dft_force \
15 --energy_key=dft_energy \
16 --stress_key=dft_stress \
17 --lr=0.0005 \
18 --scaling="rms_forces_scaling" \
19 --batch_size=4 \
20 --max_num_epochs=150 \
21 --ema \
22 --ema_decay=0.99 \
23 --weight_decay=1e-6 \
24 --amsgrad \
25 --default_dtype="float64" \
26 --clip_grad=10 \
27 --device=cuda \
28 --seed=\$SEED \
29 --num_samples_pt=500 \
30 --swa_energy_weight=100.0 \
31 --swa_forces_weight=10.0 \
32 --swa_stress_weight=1.0 \
33 --swa \
34 --swa_lr=5e-4 \
35 --E0s="{3:-0.29745415, 15:-1.88719893, 16:-1.08090657, 32:-0.77926738}" \
Listing 4: Fine-tuning template for the MACE-MP-0 model.
Based on our experience, we suggest using at least $L=1$ (i.e. medium) models for a balance between accuracy and efficiency. The following recommendations are in general helpful on stability and accuracy of the resulting model from fine-tuning:
  • Isolated atom energies (E0s): It is very important to compute your own E0s, which specify the isolated atom energy as a dictionary in your DFT calculation rather by estimating it from least square regression with E0s="average". When computing your E0s, use spin polarized calculations. You may also use E0s=”foundation” to use the same E0s when using exactly the same DFT settings and software as Materials Project.
  • Learning rate (lr): If the training or validation loss stagnates or diverges, consider modifying the learning rate from recommended range for more stable convergence.
  • Loss weights: Depending on the target properties, modify the ratio between energy, forces and stress weight in fine-tuning within the recommended range.
  • Cutoff radius (r_max): Do not modify the pre-defined cutoff radius that the foundation model used.
  • Batch size: A larger batch size generally improves training stability but requires more GPU memory. Batch size larger than 20 is not recommended since a larger batch size, in general, tend to deteriorate the generalizibility of machine learning models.
  • EMA and SWA: Enable exponential moving average (ema) and stochastic weight averaging (swa) for smoother convergence and better generalization. If training dynamics appear unstable, set swa to 0.99999. The model with best validation error from either pre-swa or post-swa based (i.e. stageone and stagetwo in Section II.4) on validation loss are outputted to be further verified on other properties.
  • Number of samples in the pretrained head (num_samples_pt): A larger value helps preserve more information and improves stability in molecular dynamics. However, there is a trade-off: an overly large value may reduce the achievable accuracy.
We recommend monitoring convergence behavior (cf. Fig. 2) and validation accuracy throughout the fine-tuning process. Evaluating model performance on unseen configurations is crucial and will be discussed in the next subsection.
II.4 Model Validation and Evaluation
After completing the fine-tuning process, several output files are generated by the training script. Among them, the most critical is the log file, which contains all essential information needed to reproduce the results. This includes the selected hyperparameters, loss trajectories over all training epochs, and final test errors (typically reported in terms of RMSE). Toward the end of the log, a summary table presents test errors for energies, forces, and—if available—stresses.
Additionally, the script automatically generates diagnostic plots that visualize the training and validation loss curves, as well as scatter plots comparing predicted and reference energies and forces. These visualizations provide a qualitative assessment of model performance. In particular, smooth loss convergence and tightly clustered scatter plots are indicative of successful fine-tuning.
As an illustration, we refer to the example in Section III.1, where such diagnostic plots are used to evaluate the performance of a fine-tuned model. Figure 2 displays the evolution of training and validation losses, along with the final energy and force scatter plots.
Figure 2: Training and validation loss curves, and fitting accuracy for energy and force predictions. The left panel corresponds to the initial training stage (before stochastic weight averaging), while the right panel shows results after SWA.
The output directory also contains several key model files:
  • MACE-FT.model: the complete fine-tuned model used for inference;
  • MACE-FT_compiled.model: a compiled version optimized for deployment;
  • MACE-FT_stageone.model and MACE-FT_stagetwo.model: intermediate models from different training stages (e.g., before and after stochastic weight averaging).
These models can be directly used in downstream atomistic simulations and analysis workflows to evaluate a variety of important material properties. These include structural and mechanical properties such as bulk modulus and defect formation and migration energies, thermodynamic properties such as density, pressure, and phonon spectra, as well as dynamical properties such as radial distribution functions and diffusion coefficients obtained from molecular dynamics simulations.
Example: Structural Relaxation with ASE.
Once fine-tuned, the model can be seamlessly integrated into the Atomic Simulation Environment (ASE) for structural optimization and property evaluation. The example below demonstrates how to perform structural relaxation on a given atomic configuration using the MACECalculator within ASE:
1from mace.calculators import MACECalculator
2ase_atoms.calc = MACECalculator("MACE-FT.model")
3
4from ase.optimize.precon import PreconLBFGS
5optimizer = PreconLBFGS(ase_atoms, variable_cell=True)
6optimizer.run(fmax=1e-5)
Listing 5: Structural relaxation using a fine-tuned MACE model
This procedure minimizes both atomic positions and cell parameters using the preconditioned L-BFGS algorithm. The same workflow can be extended to defect formation energy calculations, nudged elastic band (NEB) transition state searches, or molecular dynamics (MD) simulations.
Example: Using the Model in LAMMPS.
To use the fine-tuned model in LAMMPS, it must first be converted to a LAMMPS-compatible format using the provided utility script:
1python <mace_repo_dir>/mace/cli/create_lammps_model.py MACE-FT.model
Listing 6: Exporting the model for LAMMPS
Executing the above command generates a LAMMPS-compatible model file named MACE-FT.model-lammps.pt, which can be directly used in LAMMPS simulations via the following commands:
1pair_style mace no_domain_decomposition
2pair_coeff * * MACE-FT.model-lammps.pt elmemt1 elment2 elment3
Listing 7: Using the fine-tuned MACE potential in LAMMPS
Here, pair_style specifies the MACE interaction style without domain decomposition, and pair_coeff is used to load the compiled model file along with the corresponding atomic species. In our molecular dynamics simulations, we adopted the metal unit style and applied a canonical (NVT) ensemble to maintain the temperature at 298 K.
We refer readers to our third case study (Section III.3) for a complete example of running MD simulations in LAMMPS using a fine-tuned model. All input files and scripts are available in our public tutorial repository: https://github.com/John2021-hub/mace-ft-tutorial.git.
III EXAMPLES AND APPLICATIONS
The fine-tuned models integrate seamlessly with widely used atomistic simulation platforms such as ASE, LAMMPS and RBMD (see Appendix V.3), enabling efficient molecular dynamics and other atomistic simulations. In this section, we illustrate the effectiveness of fine-tuning using representative case studies on solid-state electrolytes, metallic stacking faults, and interfacial phenomena in low-dimensional systems. The pretrained foundation models are publicly available at https://github.com/ACEsuit/mace-mp/. In addition, we provide fine-tuning scripts, example outputs, and best practices in our tutorial repository: https://github.com/John2021-hub/mace-ft-tutorial.git.
III.1 Li10GeP2S12
In this example, we demonstrate the fine-tuning procedure on a solid-state electrolyte system, lithium germanium phosphorous sulfide (Li10GeP2S12, abbreviated as LGPS), using a dataset published on DPAISquare [73]. LGPS is a prototypical member of the argyrodite family and has attracted significant attention due to its high lithium-ion conductivity, making it a promising candidate for next-generation all-solid-state batteries. Modeling such materials requires accurate interatomic potentials capable of describing both local structural features and ionic transport mechanisms, making LGPS an ideal testbed for evaluating the transferability and fine-tuning of U-MLIPs.
The LGPS dataset published on DPAISquare [73] can be converted into the .xyz format supported by MACE, making it directly usable in the fine-tuning workflow (cf. Listing 3). Figure 3 shows the atomic structure of a typical configuration from this dataset.
Figure 3: Atomic structure of Li10GeP2S12, a solid-state electrolyte with high lithium-ion conductivity.
Table 4 summarizes the test accuracy of multiple MACE foundation models after fine-tuning. All models achieve comparable accuracy in terms of root-mean-square error (RMSE) with respect to DFT reference values. The two rows for each model correspond to results obtained before and after the application of stochastic weight averaging (SWA), where the first epoch indicates the SWA start step, and the second corresponds to the final averaged model. The consistency of test accuracy before and after SWA highlights the stability and robustness of the fine-tuning procedure. Also, the key point is that for all U-MLIPs, with the same hyperparameters, we can all get comparable accuracy, showing that the robustness of MACE-based foundation models and its fine-tuning.
Table 4: Test accuracy of different fine-tuned MACE foundation models, reported as RMSE of energy (E) and force (F), relative to DFT values. “Relative F” denotes the force RMSE normalized by the standard deviation of DFT forces. Each model is evaluated before and after SWA, with the epoch number indicating the training step.
Model (epoch) RMSE E (meV/atom) RMSE F (meV/Å) Relative F (%)
MACE-0b3 (110) 0.32 16.15 1.92
MACE-0b3 (148) 0.28 15.84 1.87
MACE-mpa (110) 0.30 15.07 1.79
MACE-mpa (146) 0.27 14.88 1.77
MACE-omat (110) 0.35 15.46 1.84
MACE-omat (149) 0.27 15.32 1.82
Figure 4 compares the test accuracy of models fine-tuned from different MACE foundation models with that of training-from-scratch, evaluated at different fractions of the training dataset (10%, 50%, and 100%). Figure 4(a) shows the RMSE in energy, while Figure 4(b) shows the RMSE in forces. Across all dataset sizes, fine-tuned models (mp-0b3, mpa-0, omat-0) consistently achieve lower errors than training-from-scratch, particularly at lower data fractions, demonstrating the data efficiency and accuracy benefits of fine-tuning. The results highlight the robustness of MACE foundation models as effective starting points for fine-tuning, leading to improved performance with reduced training data.
Figure 4: Comparison between fine-tuning and training-from-scratch in terms of data efficiency: (a) RMSE in energy (meV/atom) and (b) RMSE in forces (meV/Å).
III.2 Molybdenum
In this example, we evaluate the capability of fine-tuned U-MLIPs to accurately describe mechanical properties associated with crystalline defects in metallic systems. Molybdenum (Mo), a body-centered cubic (BCC) transition metal, is chosen as a representative case due to its well-characterized elastic properties and defect energetics, including vacancy formation and migration [105].
Crystalline defects play a key role in determining the plasticity, strength, and diffusion behavior of metals. However, accurately capturing their highly localized and stress-sensitive nature remains a challenge for many interatomic potentials. This example is therefore designed to assess whether fine-tuning improves the quantitative prediction of defect-related energies and transition states in Mo. The atomic structure of bulk Mo is shown in Figure 5. The dataset used for fine-tuning is taken from Ref. [105].
Figure 5: Atomic structure of BCC Mo containing a dislocation and the associated local lattice distortion. The defect core is highlighted in red.
Bulk validation — Table 5 compares the elastic constants ($C_{11}$, $C_{12}$, $C_{44}$), bulk modulus ($B$), and Poisson ratio ($\nu$) of molybdenum predicted by different interatomic potentials (EAM, GAP, MACE-MP-0b3, and fine-tuned MACE-MP-0) against DFT and experimental reference values. The numbers in parentheses indicate the percentage error relative to experiment. The results show that fine-tuning the MACE-MP-0 model significantly improves its accuracy across all properties, reducing the error from 45.91% to 2.58% for $C_{11}$, from 56.88% to 14.77% for $C_{44}$, and from 48.27% to 3.45% for $\nu$. While the original MACE-MP-0b3 underestimates elastic constants and bulk modulus, fine-tuning closes the gap to both DFT and experimental values, achieving accuracy comparable to or better than EAM and GAP in most properties. These results highlight the effectiveness of fine-tuning in improving the mechanical property predictions of U-MLIPs.
Table 5: Elastic constants $C_{ij}$, bulk modulus $B$, and Poisson ratio $\nu$, as obtained with the GAP, tabGAP, EAM/FS, and the NNIP potentials, compared to DFT calculations performed in this work and experimental data. Values in parentheses report the magnitude of the percentage error with respect to the experimental value.
EAM [106] GAP [107] MACE-MP-0b3 Fine-tuning DFT [105] Experiment [108]
$C_{11}$ (GPa) 465 (0.22%) 478 (3.02%) 251 (45.91%) 452 (2.58%) 459 464
$C_{12}$ (GPa) 161 (1.26%) 166 (4.40%) 189 (18.87%) 174 (9.43%) 162 159
$C_{44}$ (GPa) 109 (0%) 108 (0.92%) 47 (56.88%) 82 (14.77%) 97 109
$B$ (GPa) 263 (5.20%) 270 (8.00%) 210 (16.00%) 267 (6.80%) 262 250
$\nu$ 0.26 (10.34%) 0.26 (10.34%) 0.43 (48.27%) 0.28 (3.45%) 0.30 0.29
Generalized stacking fault energy — Figure 6 shows the generalized stacking fault energy (GSFE) profiles of Mo along the $\langle 110\rangle$ (panel (a)) and $\langle 112\rangle$ (panel (b)) directions. Table 6 provides the corresponding GSFE barrier errors relative to DFT. The results compare the MACE foundation model, two fine-tuned models (FT-Default, which employs the default loss weights, and FT, which uses an increased force loss weight of 100.0), and the DFT reference data. The foundation model clearly underestimates the GSFE in both directions, leading to lower energy barriers. After fine-tuning, the predictions become much closer to the DFT values, especially near the peak and along the entire fault path. Between the two fine-tuned models, FT performs slightly better than FT-Default, suggesting that adjusting the force weighting can further improve accuracy.
Table 6: GSFE barrier errors relative to DFT are reported for the $\langle 110\rangle$ and $\langle 121\rangle$ slip directions, comparing three interatomic potential models: Fine-Tuning, FT-Default, and MACE-MP-0b3.
Slip Directions Fine-Tuning FT-Default MACE-MP-0b3
$\langle 110\rangle$ 0.17 0.23 0.88
$\langle 121\rangle$ 0.21 0.36 0.78
Figure 6: GSFE profiles of Mo along the $\langle 110\rangle$ (panel (a)) and $\langle 112\rangle$ (panel (b)) directions. Two fine-tuned models are compared: FT-Default, which uses the default loss weights, and FT, which employs a force loss weight of 100.0.
III.3 Graphene–Water Interface
Graphene–water interfaces are prototypical solid–liquid systems with wide-ranging applications in sensing, electrochemistry, and nanofluidics. Capturing the interfacial structure and dynamics of such systems presents a significant challenge for interatomic potentials, especially in describing long-range interactions and subtle surface effects. This example illustrates the capability of fine-tuned U-MLIPs to model solid–liquid interfaces accurately, using a representative graphene–water system. The atomic structure of the graphene–water interface is illustrated in Figure 7. The system contained over 372 atoms and was simulated for 200 ps with a 1 fs timestep.
Figure 7: Atomic structure of the graphene–water interface used in this study.
Figure 8 compares molecular dynamics simulations of a graphene–water interface performed using LAMMPS-MACE with the fine-tuned model (MACE-FT-MD) and ab initio molecular dynamics (AIMD). Figure 8(a) shows the temperature evolution, with both simulations maintaining comparable temperature fluctuations around 300 K. Figure 8(b) presents the oxygen number-density profile along the $z$-axis. The first peak corresponds to the graphene–water interfacial distance, ${\rm d}{\rm gw}=0.35$ nm, which is in good agreement with the experimental value of 0.36 nm [109]. Due to the hydrophobic nature of graphene, a depletion layer forms at the interface, and ${\rm d}{\rm gw}$ is defined to quantify this interfacial gap. This distance is a critical physical quantity for evaluating the accuracy of the interaction potentials and demonstrates the effectiveness of the fine-tuned model.
Figure 8: Comparison of graphene–water interface simulations using MACE-FT-MD and AIMD. (a) Temperature evolution over time. (b) Oxygen number-density profile along the z-axis; The first peak corresponds to the graphene–water distance, ${\rm d}_{\rm gw}=0.35~{}{\rm nm}$, which is in good agreement with the experimental value of 0.36 nm [109]. The inset illustrates the definition of ${\rm d}_{\rm gw}$.
Table 7 reports the computational cost (wall-clock time in seconds) for 10,000-step molecular dynamics simulations performed with and without the cuEquivalence kernel, across systems of varying sizes (128 to 2900 atoms). For each configuration, simulations using cuEquivalence (“With”) show consistently lower elapsed times compared to runs without it (“Without”), especially for larger systems. For instance, at 2900 atoms, cuEquivalence reduces the runtime from 5084 seconds to 2619 seconds. The scaling factor improves from 0.76 (without cuEquivalence) to 0.41 (with cuEquivalence), indicating better parallel efficiency with the kernel enabled.
Table 7: Computational cost (wall-clock time in seconds) for 10000-step molecular dynamics simulations with and without the use of the cuEquivalence kernel. The first column indicates the total number of atoms in the system, while the remaining columns report the elapsed time for each configuration. “With” denotes simulations utilizing the cuEquivalence kernel, and “Without” corresponds to runs without it.
Atoms 128 372 1038 2900 Scaling
With (s) 710 800 1126 2619 0.41
Without (s) 658 1008 2031 5084 0.76
IV Summary and Perspective
This tutorial provides a practical guide to fine-tuning universal machine-learned interatomic potentials (U-MLIPs), using the MACE-MP-0 model as a representative example. We cover the end-to-end process—from dataset preparation and hyperparameter selection to model training and application—with the goal of helping researchers efficiently adapt pre-trained models to their systems of interest. The effectiveness of fine-tuning is demonstrated through representative examples, including solid-state electrolytes, stacking fault defects in metals, and surface interactions in low-dimensional materials. To support reproducibility, all code and data are made publicly available, along with scripts for running molecular dynamics simulations using the RBMD platform.
While this tutorial focuses on MACE-MP-0, the strategies discussed are readily transferable to other U-MLIP frameworks. We hope that this work will help lower the barrier to entry for applying fine-tuned MLIPs in atomistic simulations, and encourage their broader use in both academic and industrial settings.
That said, caution is required when fine-tuning or applying MLIPs. These models approximate the potential energy surface purely from data and do not inherently incorporate physical laws or constraints. Their accuracy is ultimately determined by the coverage and quality of the training dataset. We point out several important considerations and open directions for future development of MLIPs.
Long-Range Interactions.
Most MLIPs rely on a local decomposition of the total energy, in which each atom’s contribution is determined by its local environment within a finite cutoff radius (typically 5–10 Å). While this approximation is effective in condensed-phase systems where long-range interactions are screened, it becomes inadequate in systems such as ionic solids, electrolytes, and interfacial structures, where electrostatic and dispersion interactions extend well beyond the cutoff. In these cases, long-range effects may be misinterpreted as noise during training, ultimately degrading model accuracy. To address this issue, several extensions have been proposed [110, 111, 112, 113, 114], including environment-dependent charge models, DFT-D-type dispersion corrections, and hybrid approaches that integrate ML models with explicit physics-based components. Continued progress in this direction is expected, as long-range interactions remain a key challenge and an active area of research in the development of accurate and transferable MLIPs.
Multiple Electronic States.
Standard MLIPs are designed to model a single potential energy surface, typically the ground state under the Born–Oppenheimer approximation [115]. However, many important phenomena—including spin transitions, charge transfer, and photocatalysis—require access to excited-state dynamics. Representing multiple electronic states remains a challenge due to the need to model sharp energy differences and nonadiabatic couplings. Some approaches train separate ML models for different states [116, 117, 118], while others explore multi-output architectures [119]. Recent work [120] has also focused on spin-dependent descriptors and graph neural networks with embedded electronic information, which show promise in this direction.
Uncertainty Quantification and Active Fine-Tuning.
Uncertainty quantification (UQ) is essential for evaluating model reliability and guiding data selection in active learning and fine-tuning. Techniques such as ensemble variance, dropout inference, and Bayesian approximations help identify regions of high predictive uncertainty, enabling targeted sampling and noise reduction. In fine-tuning, UQ reduces data redundancy and ensures that added configurations meaningfully enhance model performance. A central challenge is the automated construction of fine-tuning datasets informed by uncertainty, task relevance, or physical constraints. Once such datasets can be systematically generated, selecting or adapting models becomes significantly more efficient. Developing robust pipelines for uncertainty-aware data generation is thus key to scalable, task-adaptive fine-tuning [121].
Building Intermediate-Sized MLIPs for Specific Tasks.
Although U-MLIPs demonstrate excellent generalization across a wide range of materials systems, they may not be optimal for small- to medium-sized systems or domain-specific tasks. In such cases, it may be more effective to either fine-tune a compact pre-trained model or train a lightweight architecture from scratch on a focused dataset [122, 123]. This approach offers a favorable trade-off between interpretability, computational efficiency, and predictive accuracy—particularly when computational resources are limited or when real-time inference is required for large-scale simulations. Lightweight variants of ACE and other descriptor-based models remain valuable tools for such targeted applications.
The Need for Comprehensive Benchmarking Platforms for MLIPs.
As MLIPs continue to evolve into foundation models with broad applicability, there is an increasing need for benchmarking platforms that go beyond traditional accuracy metrics such as RMSE. Evaluating an MLIP’s true usefulness requires not only numerical accuracy, but also physical fidelity, generalization ability, and reliability in practical simulations. Conventional benchmarks often fail to capture these dimensions, making it difficult to assess whether a model is suitable for deployment in real-world applications. To address this gap, recent efforts have begun to develop more comprehensive evaluation frameworks that emphasize model robustness, transferability, and physics-informed behavior. One notable example is MLIP Arena [124], a newly proposed platform that offers architecture-agnostic, task-oriented benchmarking of open-source, open-weight MLIPs. It represents a promising step toward standardized, holistic assessment of model performance in complex materials modeling tasks.
Post-Training Strategies.
After pre-training a foundation model, post-training serves as a crucial step to further adapt and optimize the model for specific downstream applications. One effective post-training technique is model distillation, in which a simpler model (the “student”) is trained to replicate the behavior and predictive performance of a larger, more complex model (the “teacher”—in this case, U-MLIPs). This is typically achieved by using the teacher’s outputs or internal representations as supervisory signals for training the student [125]. Recent work [126] demonstrates that matching the Hessians of energy predictions between the teacher and student enables effective distillation. In particular, large-scale foundation models can be distilled into compact, task-specific MLIPs tailored to a particular chemical subdomain, achieving up to 50$\times$ speedup in inference time compared to the original foundation model. Future investigations on this topic will be conducted using the RBMD platform.
V Appendices
This appendix provides supplementary materials to complement the main content of the tutorial. Section V.1 presents a concise overview of existing U-MLIPs, offering historical context and highlighting key differences among representative models. Section V.2 provides additional technical details on the MACE architecture for readers seeking a deeper understanding of its design and capabilities. Finally, Section V.3 compiles a list of useful code repositories and publicly available datasets to support reproducibility and facilitate the practical application of the methods described in this work.
V.1 Review of U-MLIPs
While recent advances in universal machine-learned interatomic potentials (U-MLIPs) are largely driven by modern machine learning techniques, their conceptual roots can be traced to earlier semi-universal, physics-based force fields (SU-PBPs). Early examples of SU-PBPs include the work of Weiner et al. [127] in 1981, which laid the foundation for general-purpose molecular potentials. This line of development continued with the introduction of widely used models such as the Universal Force Field (UFF) by Rappé et al. [128] and the AMBER family of force fields [127, 129], both of which have seen broad application in molecular simulations, particularly in fields like pharmaceutical research.
Despite their shared aim of general applicability, SU-PBPs and U-MLIPs differ significantly in methodology and maturity. While SU-PBPs are based on empirical and physically motivated functional forms, U-MLIPs are trained directly on quantum mechanical data using flexible machine-learned architectures. A pioneering example of a semi-universal MLIP is the ANI (Accurate NeurAl network engINe for Molecular Energies) model, introduced by Smith et al.[11, 130]. Originally trained on organic molecules containing C, H, O, and N, ANI was later expanded to incorporate S, Cl, and F, thereby achieving coverage of approximately 90% of drug-like chemical space[131]. Functionally, ANI can be seen as a machine-learned counterpart to traditional force fields like AMBER.
Further developments have extended the scope and accuracy of ANI-type models. For instance, AIMNet2 [132] includes 14 elements and introduces corrections for long-range interactions, such as dispersion and electrostatics, enabling improved modeling of charged systems and diverse bonding environments. Similarly, ANI-1xnr [133] adapts the ANI framework for use in condensed-phase systems, including liquids and reactive environments.
More recently, the MACE-OFF23 model [134] has leveraged the MACE message-passing architecture to construct a short-range, transferable potential for the ten most prevalent elements in organic chemistry (H, C, N, O, F, P, S, Cl, Br, I). Despite not explicitly modeling long-range interactions, MACE-OFF23 has shown performance improvements over ANI across several benchmarks, highlighting the progress in developing efficient and accurate ML-based universal potentials.
The first truly universal MLIPs with broad elemental coverage across the periodic table emerged around 2022 [27, 135]. Since then, many U-MLIPs have been proposed, including:
  • M3GNet [27]: a 3-body materials graph network developed by Chen and Ong;
  • CHGNet [23]: the Crystal Hamiltonian Graph Neural Network;
  • ALIGNN-FF [26]: a unified atomistic force field based on line graph neural networks;
  • PFP [33]: a tensorial message-passing network developed by Takamoto et al., also released commercially in Matlantis;
  • GNoME [24]: based on NequIP [8], trained on $\sim$80 million in-house DFT calculations;
  • SevenNet-0 [34]: optimized for large-scale simulation with better parallel scaling;
  • MACE-MP-0 [6]: a U-MLIP based on the MACE architecture [6], trained on the same public data as CHGNet and SevenNet, demonstrating strong performance across aqueous systems, catalysis, and MOFs;
  • GPTFF [136]: a GNN-transformer hybrid force field;
  • MatterSim [35]: trained by Microsoft on a large dataset of $\sim$17 million structures covering broad thermodynamic conditions;
  • Orb [137]: a high-speed CUDA-optimized model offering 3–6$\times$ speedup over PyTorch-based MACE;
  • Equiformer V2-OMAT24 [36]: currently ranked highest on the MatBench leaderboard, trained on $\sim$118 million structures.
There are several recent studies have systematically benchmarked U-MLIPs [138, 139, 41, 140]. These evaluations underscore the importance of assessing model performance not only in terms of energy and force errors, but also with respect to physically meaningful properties such as surface energies, elastic moduli, and defect energetics—quantities that are often more relevant for practical applications. However, constructing diverse and high-quality test sets for such evaluations remains an open challenge. A key concern is the risk of overfitting when test datasets become implicit targets during model optimization, which may compromise generalization to unrelated properties or systems. Moreover, it is essential to recognize that low errors in energy and force predictions do not necessarily ensure reliable performance on derived physical properties. This limitation is inherent to MLIPs in general and is not specific to U-MLIPs.
V.2 MACE Architecture
In this subsection, we introduce the main idea behind the construction of the MACE architecture. We provide only a brief overview here and refer interested readers to [6] for more detailed constructional insights.
First, a graph is defined by connecting two nodes (atoms) with an edge if they are within each other’s local environment. The local environment, $\mathcal{N}(i)$, consists of all atoms $j$ surrounding a central atom $i$ such that $\|\mathbf{r}_{ij}\|\leq r_{\text{cut}}$, where $\mathbf{r}_{ij}$ is the vector from atom $i$ to atom $j$, and $r_{\text{cut}}$ is a cutoff hyperparameter. The feature vector of node $i$ is denoted by $\mathbf{h}_{i}^{(t)}$, expressed in the spherical harmonic basis, with indices $l$ and $m$. The superscript $t$ represents the iteration step (analogous to the “layers” in graph neural networks’ message-passing). These node features, $\mathbf{h}_{i}^{(t)}$, reflect the chemical environment of the atoms.
It is worth noting that all MACE models consist of only two layers. Due to the specific construction, MACE can be regarded as a higher-order message-passing neural network. This enables it to achieve comparable expressiveness with just two layers, which is a key strength of the MACE model.
The node features on 0-th layer denoted as $\mathbf{h}_{i}^{(0)}$ are initialized as a (learnable) embedding of the chemical elements with atomic numbers $z_{i}$ into $k$ channels:
$h_{i,k00}^{(0)}=\sum_{z}W_{kz}\delta_{zz_{i}}.$ (1)
This type of mapping has been widely applied to graph neural networks [7, 141] and other models [142, 38].
Next, for each atom, the features of its neighbors are combined with the interatomic displacement vectors, $\mathbf{r}_{ij}$, to form the one-particle basis $\phi_{ij,knl_{1}l_{2}m_{2}}^{(t)}$. The construction of MACE employs the idea from the ACE architecture, which is a natural many-body representation for capturing the symmetry of many-body particle systems. The radial distance $r_{ij}$ is used as an input into a learnable radial function $R(r_{ij})$, with several outputs that correspond to different ways in which the displacement vector and node features can be combined while maintaining equivariance [143]:
$\phi_{ij,knl_{1}l_{2}m_{3}}^{(t)}=\sum_{l_{1}m_{1}l_{2}m_{2}}$ $C_{\eta_{1},l_{1}m_{1},l_{2}m_{2}}^{l_{3}m_{3}}R_{knl_{1}l_{2}l_{3}}^{(t)}(r_{ij})$
$Y_{l_{1}m_{1}}^{m_{1}}(\hat{r}_{ij})h_{j,kl_{2}m_{2}}^{(t)},$ (2)
where $Y_{l}^{m}$ are the spherical harmonics, and $C_{\eta_{1},l_{1}m_{1},l_{2}m_{2}}^{l_{3}m_{3}}$ denotes the Clebsch-Gordan coefficients. There are multiple ways of constructing an equivariant combination with a given symmetry $(l_{3},m_{3})$, and these multiplicities are enumerated by the path index $\eta_{1}$ [37].
The one-particle basis $\phi$ is summed over the neighborhood and expanded as a linear combination of $k$ channels with learnable weights to form the permutation-invariant atomic basis $A_{i}$:
$A_{i,kl3m3}^{(t)}=\sum_{\tilde{k},\eta_{1}}W_{\tilde{k}\tilde{k}k\eta_{1}l_{3}}\sum_{j\in\mathcal{N}(i)}\phi_{ij,\tilde{k}\eta_{1}l_{3}m_{3}}^{(t)}.$ (3)
Higher-order (many-body) symmetric features are created on each atom by taking products of the atomic basis, $A$, with itself $\nu$ times, resulting in the “product basis”. The product basis is then contracted with the generalised Clebsch-Gordan coefficients $C_{\eta_{\nu}lm}^{LM}$ to obtain the equivariant higher-order basis, $B_{i}$ [37]:
$B_{i,\eta_{\nu}kLM}^{(t),\nu}=\sum_{lm}C_{\eta_{\nu}lm}^{LM}\prod_{\xi=1}^{\nu}A_{i,kl_{\xi}m_{\xi}}^{(t)},$ (4)
where bold $lm$ denotes the $\nu$-tuple of $l$ and $m$ values and similarly to Equation (2), $\eta_{\nu}$ enumerates the number of possible couplings to create the features with equivariance $LM$.
Finally, a “message” $m_{i}$ is created on each atom as a learnable linear combination of the equivariant many-body features:
$m_{i,kLM}^{(t)}=\sum_{\nu}\sum_{\eta_{\nu}}W_{zi\eta_{\nu}kLM}^{(t),\nu}B_{i,\eta_{\nu}kLM}^{(t),\nu}.$ (5)
The recursive update of the node features $(t:0\rightarrow 2)$ is obtained by adding the message to the atoms’ features from the previous iteration, with weights that depend on the chemical element ($z_{i}$) that are also responsible for the mixing of the chemical embedding $k$ channels:
$h_{i,kLM}^{(t+1)}=\sum_{\tilde{k}}W_{kL,\tilde{k}}^{(t)}m_{i,\tilde{k}LM}^{(t)}+\sum_{\tilde{k}}W_{kziL,\tilde{k}}^{(t)}h_{i,\tilde{k}LM}^{(t)}.$ (6)
The site energy is a sum of read-out functions applied to node features from the first and second layers. The read-out function is defined as a linear combination of rotationally invariant node features for the first layer, and as a multi-layer perceptron (MLP) for the second layer.
$E_{i}=\sum_{t=1}^{2}E_{i}^{(t)}=\sum_{t=1}^{2}\mathcal{R}^{(t)}\left(h_{i}^{(t)}\right),$ (7)
where
$\mathcal{R}^{(t)}\left(h_{i}^{(t)}\right)=\begin{cases}\sum_{k}W_{k}^{(t)}h_{i,k00}^{(t)}&\text{for }t=1\\ \text{MLP}\left(\left\{h_{i,k00}^{(t)}\right\}_{k}\right)&\text{for }t=2\end{cases}.$ (8)
The forces and stresses on the atoms are calculated by taking analytical derivatives of the total potential energy with respect to the positions of the atoms.
V.3 Codes and Data Availability
To support reproducibility, extensibility, and practical adoption of MACE-based workflows, Table 8 summarizes key publicly available GitHub repositories and data resources. These include official codebases for model construction and training, fine-tuning utilities, simulation interfaces (e.g., LAMMPS [16], ASE [15], OpenMM [144]), and model conversion or acceleration tools. The collection also contains pretrained foundation models and benchmark datasets. Together, these components provide a comprehensive ecosystem for developing and deploying MACE across diverse material systems.
Table 8: Key repositories and data resources supporting the development, deployment, and application of MACE-based models. The table includes core codebases, pretrained models, tutorials, simulation interfaces, kernel and dispersion extensions, and benchmark datasets.
Repository / Resource Purpose / Description Link / Owner
mace Official MACE codebase for model construction, training, evaluation, and inference. ACEsuit/mace
MACE-MP-0 Pretrained MACE-MP-0 foundation model checkpoints, logs, and metadata. ACEsuit/mace-foundations
mace-off MACE-OFF23 pretrained transferable organic force fields for molecular systems. ACEsuit/mace-off
mace-ft-tutorial Fine-tuning tutorial with scripts and examples for adapting pretrained MACE models. https://github.com/John2021-hub/mace-ft-tutorial
mace-tutorials-csc A tutorial for training MACE models. ilyes319/mace-tutorials-csc
mace-jax JAX-based reimplementation of MACE for TPU/GPU acceleration and efficient autodiff. ACEsuit/mace-jax
mace-layer MACE layer for neural operator applications on 3D geometric point clouds. ACEsuit/mace-layer
mace-lammps LAMMPS interface for deploying MACE potentials in molecular dynamics simulations. ACEsuit/mace-lammps
mace-md MD wrapper supporting MACE potentials, especially with OpenMM backend. jharrymoore/mace-md
RBMD Random-batch MD framework for large-scale atomistic simulations using MACE. https://www.randbatch.com/rbmd
torch-dftd PyTorch-based DFT-D3 module for dispersion correction, usable via ASE with MACE. CheukHinHoJerry/torch-dftd
cuEquivariance CUDA-accelerated kernel library for equivariant tensor operations in MACE. NVIDIA/cuEquivariance
cuda-mace Custom CUDA implementation of MACE for high-performance inference. nickjbrowning/cuda_mace
Materials Project (MP) data Dataset used to train MACE-MP-0 and CHGNet; includes DFT-calculated material properties. materialsproject.org